In a survey of 120 grade 5 pupils, 50 do not like to watch Pinocchio. Find p and q where p was the proportion of children who like to watch the film.


1
Expert's answer
2022-02-02T08:39:14-0500

Solution:

"n=120"

p was the proportion of children who like to watch the film.

And 50 do not like to watch Pinocchio.

So, (120-50=) 70 like to watch Pinocchio.

"p=\dfrac{70}{120}=\dfrac{7}{12}
\\ q=\dfrac{50}{120}=\dfrac{5}{12}"